### v3

# RevPRAG: Revealing Poisoning Attacks in Retrieval-Augmented Generation through LLM Activation Analysis

Abstract
Retrieval-Augmented Generation (RAG) enriches the input to LLMs by retrieving information from the relevant knowledge database, enabling them to produce responses that are more accurate and contextually appropriate. It is worth noting that the knowledge database, being sourced from publicly available channels such as Wikipedia, inevitably introduces a new attack surface. RAG poisoning attack involves injecting malicious texts into the knowledge database, ultimately leading to the generation of the attacker’s target response (also called poisoned response). However, there are currently limited methods available for detecting such poisoning attacks. We aim to bridge the gap in this work by introducing RevPRAG, a flexible and automated detection pipeline that leverages the activations of LLMs for poisoned response detection. Our investigation uncovers distinct patterns in LLMs’ activations when generating poisoned responses versus correct responses. Our results on multiple benchmarks and RAG architectures show our approach can achieve a 98% true positive rate, while maintaining a false positive rate close to 1%.
RevPRAG: Revealing Poisoning Attacks in Retrieval-Augmented Generation through LLM Activation Analysis
Xue Tan1, Hao Luan1, Mingyu Luo1, Xiaoyan Sun2, Ping Chen3, Jun Dai2 1School of Computer Science, Fudan University, Shanghai, China, 2Department of Computer Science, Worcester Polytechnic Institute, MA, USA, 3Institute of Big Data, Fudan University, Shanghai, China, Correspondence: xsun7@wpi.edu pchen@fudan.edu.cn jdai@wpi.edu
1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) has emerged as an effective solution that leverages retrievers to incorporate external databases, enriching the knowledge of LLMs and ultimately enabling the generation of up-to-date and accurate responses. RAG comprises three components: knowledge database, retriever, and LLM. Fig. 1 visualizes an example of RAG. The knowledge database consists of a large amount of texts collected from sources such as latest Wikipedia entries Thakur et al. (2021), new articles Soboroff et al. (2018) and financial documents Loukas et al. (2023). The retriever is primarily responsible for retrieving the texts that are most related to the user’s query from the knowledge database. These texts will later be fed to LLM as a part of the prompt to generate responses (e.g., “Everest") for users’ queries (e.g., “What is the name of the highest mountain?"). Due to RAG’s powerful knowledge integration capabilities, it has demonstrated impressive performance across a range of QA-like knowledge-intensive tasks Lazaridou et al. (2022); Jeong et al. (2024).
RAG poisoning refers to the act of injecting malicious or misleading content into the knowledge database, contaminating the retrieved texts in RAG and ultimately leading it to produce the attacker’s desired response (e.g., the target answer could be “Fuji" when the target question is “What is the name of the highest mountain?"). This attack leverages the dependency between LLMs and the knowledge database, transforming the database into a new attack surface to facilitate poisoning. PoisonedRAG  Zou et al. (2024) demonstrates the feasibility of RAG poisoning by injecting a small amount of maliciously crafted texts into the knowledge database utilized by RAG. The rise of such attacks has drawn significant attention to the necessity of designing robust and resilient RAG systems. For example, INSTRUCTRAG Wei et al. (2024) utilizes LLMs to analyze how to extract correct answers from noisy retrieved documents; RobustRAG Xiang et al. (2024) introduces multiple LLMs to generate answers from the retrieved texts, and then aggregates the responses. However, the aforementioned defense methods necessitate the integration of additional large models, incurring considerable overheads. Meanwhile, it is difficult to promptly assess whether the current response of RAG is trustworthy or not.
Figure 1: Visualization of RAG.
In our work, we shift our focus to leverage the intrinsic properties of LLMs for detecting RAG poisoning, rather than relying on external models. Our view is that if we can accurately determine whether a RAG’s response is correct or poisoned, we can effectively thwart RAG poisoning attacks. We attempt to observe LLM’s answer generation process to determine whether the response is compromised or not. It is worth noting that our focus is not on detecting malicious inputs to LLMs, as we consider the consequences of malicious responses to be far more detrimental and indicative of an attack. The growing body of research on using activations to explain and control LLM behavior Ferrando et al. (2024); He et al. (2024) provides us inspiration. Specifically, we empirically analyze the activations of the final token in the input sequence across all layers of the LLM. Our findings demonstrate that it is highly feasible to differentiate between these two types of responses by comparing the activations of the LLM when generating correct responses versus poisoned ones. Based on this, we propose a systematic and automated detection pipeline, namely RevPRAG, which consists of three key components: poisoned data collection, LLM activation collection and preprocessing, and the detection model design. It is important to note that this detection method will not alter the RAG workflow or weaken its performance, thereby offering superior adversarial robustness compared to methods that rely solely on filtering retrieved texts.
To evaluate our approach, we systematically demonstrate the effectiveness of RevPRAG across various LLM architectures, including GPT2-XL-1.5B, Llama2-7B, Mistral-7B, Llama3-8B, and Llama2-13B. RevPRAG performs consistently well, achieving over 98% true positive rate across different datasets.
Our contributions can be summarized as follows:
  1. 1.
    We uncover distinct patterns in LLMs’ activations when RAG generates correct responses versus poisoned ones.
  2. 2.
    We introduce RevPRAG, a novel and automated pipeline for detecting whether a RAG’s response is poisoned or not. To address emerging RAG poisoning attacks, RevPRAG allows new datasets to be constructed accordingly for training the model, enabling effective detection of new threats.
  3. 3.
    Our model has been empirically validated across various LLM architectures and retrievers, demonstrating over 98% accuracy on our custom-collected detection dataset.
2 Background and Related Work
2.1 Retrieval Augmented Generation
RAG comprises three components: knowledge database, retriever, and LLM. As illustrated in Fig. 1, RAG consists of two main steps: retrieval step and generation step. In the retrieval step, the retriever acquires the top $k$ most relevant pieces of knowledge for the query $q$. First, we employ two encoders, $E_{q}$ and $E_{p}$, which can either be identical or radically different. Encoder $E_{q}$ is responsible for transforming the user’s query $q$ into an embedding vector $E_{q}(q)$, while encoder $E_{p}$ is designed to convert all the information $p_{i}$ in the knowledge database into embedding vectors $E_{p}(p_{i})$. For each $E_{p}(p_{i})$, the similarity with the query $E_{q}(q)$ is computed using $\text{sim}(E_{q}(q),E_{p}(p_{i}))$, where $\text{sim}(\cdot,\cdot)$ quantifies the similarity between two embedding vectors, such as cosine similarity or the dot product. Finally, the top $k$ most relevant pieces are selected as the external knowledge $\mathcal{C}_{q}$ for the query $q$. The generation step is to generating a response $\text{LLM}(q,\mathcal{C}_{q})$ based on the query $q$ and the relevant information $\mathcal{C}_{q}$. First, we combine the query $q$ and the external knowledge $\mathcal{C}_{q}$ using a standard prompt (see Fig. 5 for the complete prompt). Taking advantage of such a prompt, the LLM generates an answer $\text{LLM}(q,\mathcal{C}_{q})$ to the query $q$. Therefore, RAG is a significant accomplishment, as it addresses the limitations of LLMs in acquiring up-to-date and domain-specific information.
2.2 Retrieval Corruption Attack
Due to the growing attention on RAG, attacks on RAG have also been widely studied. RAG can improperly generate answers that are severely impacted or compromised once the knowledge database is contaminated Zou et al. (2024); Xue et al. (2024); Jiao et al. (2024). Specifically, an attacker can inject a small amount of malicious information onto a website, which is then retrieved by RAG Greshake et al. (2023). PoisonedRAG Zou et al. (2024) injects malicious text into the knowledge database, and formalizes the knowledge poisoning attack as an optimization problem, thereby enabling the LLM to generate target responses selected by the attacker. GARAG Cho et al. (2024) was introduced to provide low-level perturbations to RAG. PRCAP Zhong et al. (2023) injects adversarial samples into the knowledge database, where these samples are generated by perturbing discrete tokens to enhance their similarity with a set of training queries. These methods have yielded striking attack results, and in our work, we have selected several state-of-the-art attack methods as our base attacks on RAG.
2.3 The Robustness of RAG
Efforts have been made to develop defenses in response to poisoning attacks and noise-induced disruptions. RobustRAG Xiang et al. (2024) mitigates the impact of poisoned texts through a voting mechanism, while INSTRUCTRAG Wei et al. (2024) explicitly learns the denoising process to address poisoned and irrelevant information. Other approaches to enhance robustness include prompt design Cho et al. (2023); Press et al. (2023), plug-in models Baek et al. (2023), and specialized models Yoran et al. (2023); Asai et al. (2023). However, these methods may, on one hand, rely on additional LLMs, leading to significant overhead. On the other hand, they primarily focus on defense mechanisms before the LLM generates a response, making it challenging for these existing approaches to detect poisoning attacks in real-time while the LLM is generating the response Athalye et al. (2018); Bryniarski et al. (2021); Carlini and Wagner (2017); Carlini (2023); Tramer et al. (2020). In this work, to defend against RAG attacks, we present a detection mechanism that can promptly capture key information during the model’s response generation process and assess whether the current response is trustworthy or not.
3 Preliminary
3.1 Threat Model
Attacker’s goal. We assume that the attacker preselects a target question set $Q$, consisting of $q_{1},q_{2},\cdots,q_{n}$, and the corresponding target answer set $A$, represented as $a_{1},a_{2},\cdots,a_{n}$. The attacker’s goal is to compromise the RAG system by contaminating the retrieval texts, thereby manipulating the LLM to generate the target response $a_{i}$ for each query $q_{i}$. For example, the attacker’s target question $q_{i}$ is “What is the name of the highest mountain?", with the target answer being “Fuji".
Attacker’s capabilities. We assume that an attacker can inject $m$ poisoned texts $P$ for each target question $q_{i}$, represented as $p_{i}^{1},p_{i}^{2},...,p_{i}^{m}$. The attacker does not possess knowledge of the LLM utilized by the RAG, but has white-box access to the RAG retriever. This assumption is reasonable, as many retrievers are openly accessible on platforms like HuggingFace. The poisoned texts can be integrated into the RAG’s knowledge database through two ways: the attacker publishing the malicious content on open platforms like Wikipedia, or utilizing data collection agencies to disseminate the poisoned texts.
Figure 2: t-SNE visualizations of activations for correct and poisoned responses.
Figure 3: The workflow of RevPRAG.
3.2 Rationale
The activations of LLMs represent input data at varying layers of abstraction, enabling the model to progressively extract high-level semantic information from low-level features. The extensive information encapsulated in these activations comprehensively reflects the entire decision-making process of the LLM. The activations has been applied to factual verification of the output content  He et al. (2024) and detection of task drift Abdelnabi et al. (2024). Due to the fact that LLM produces different activations when generating varying responses, we hypothesize that LLM will also exhibit distinct activations when generating poisoned responses compared to correct ones. Fig. 2 presents the visualizations of activations for correct and poisoned responses using t-SNE (t-Distributed Stochastic Neighbor Embedding). It visualizes the mean activations across all layers for two LLMs, Mistral-7B and Llama2-7B, on the Natural Questions dataset. This clearly demonstrates the distinguishability between the two types of responses, to some extent, supports our conjecture.
Figure 4: An instance of using RevPRAG.
4 Methodology
4.1 Approach Overview
As illustrated in Fig. 3, we introduce RevPRAG, a pipeline designed to leverage LLM activations for detecting knowledge poisoning attacks in RAG systems. It contains three major modules: poisoning data collection, activation collection and preprocessing, and RevPRAG detection model design. Fig. 4 demonstrates a practical application of RevPRAG for verifying the poisoning status of LLM outputs. Given a user prompt such as “What is the name of the highest mountain?”, the LLM will provide a response. Meanwhile the activations generated by the LLM will be collected and analyzed in RevPRAG. If the model classify the activations as poisoned behavior, it will flag the corresponding response (such as "Fuji") as a poisoned response. Otherwise, it will confirm the response (e.g. "Everest") as the correct answer.
4.2 Poisoning Data Collection
Our method seeks to extract the LLM’s activations that capture the model’s generation of a specific poisoned response triggered by receiving poisoned texts at a given point in time. Therefore, we first need to implement poisoning attacks on RAG that can mislead the LLM into generating target poisoned responses. There are three components in RAG: knowledge database, retriever, and LLM. In order to successfully carry out a poisoning attack on RAG and compel the LLM to generate the targeted poisoned response, the initial step is to craft a sufficient amount of poisoned texts and inject them into the knowledge database. In this paper, in order to create effective poisoned texts for our primary focus on detecting poisoning attacks, we employ three state-of-the-art strategies (i.e., PoisonedRAG Zou et al. (2024), GARAG Cho et al. (2024), and PAPRAG Zhong et al. (2023)) for generating poisoned texts and increasing the similarity between the poisoned texts and the queries, to raise the likelihood that the poisoned texts would be selected by the retriever. The retrieved texts, along with the question, are then used to construct a new prompt for the LLM to generate the answer. The prompt Zou et al. (2024), as shown in Fig. 5, is employed to achieve this objective.
Figure 5: The prompt used in RAG to make an LLM generate an answer based on the retrieved texts.
4.3 Activation Collection and Processing
For an LLM input sequence $X=\left(x_{1},x_{2},\cdots,x_{n}\right)$, we extract the activations $Act_{n}$ for the last token $x_{n}$ in the input across all layers in the LLM as a summary of the context. The activations $Act_{n}$ contain the inner representations of the LLM’s knowledge related to the input. When the LLM generates a response based on a question, it traverses through all layers, retrieving knowledge relevant to the input to produce an answer Meng et al. (2023). As mentioned earlier, there is a significant difference between the LLM activations for the poisoned responses and the activations for correct responses.
We introduce normalization of the activations for effective integration into the training process. We calculate the mean $\mu$ and standard deviation $\sigma$ of the dataset. Then, we use the obtained $\mu$ and $\sigma$ to normalize the activations with the formula: $Act^{nor}_{n}=\left(Act_{n}-\mu\right)/\sigma$. This standardization improves training efficiency by scaling activations uniformly, preventing bias towards larger features. It also ensures smoother optimization, alleviates gradient-related issues, and enhances the performance of algorithms that rely on distance metrics.
4.4 RevPRAG Model Design
After collecting and preprocessing the dataset of activations, we design a probing mechanism based on the dataset. Inspired by few-shot learning and Siamese networks, our proposed RevPRAG model is adept at distinguishing correct answers from poisoned ones and demonstrates strong generalizability from limited data.
In order to efficiently capture the relationships between and within different layers of the LLM, we utilize Convolutional Neural Networks (CNNs) with the ResNet18 architecture He et al. (2016). We use triplet networks which share the same architecture and weights to learn embeddings of the tasks as shown in Fig. 3. During training, we employ the triplet margin loss Schroff et al. (2015), a commonly used approach for tasks where it is difficult to distinguish similar instances. Triplet margin loss is a loss function used in training neural networks for tasks such as face recognition or object classification. At the same time, triplet margin loss is also widely used in few-shot learning scenarios. When model encounters out-of-distribution data, it can quickly adapt with a small amount of support data, without the need to retrain the entire model. The goal of this loss function is to learn a similarity metric within a high-dimensional embedding space (also known as feature space), where representations of similar objects (e.g., images of the same person) are close to each other, while representations of dissimilar objects are farther apart. This powerful similarity metric provided by triplet margin loss is particularly suitable for distinguishing LLM activations, enabling RevPRAG to effectively differentiate activation differences caused by poisoning attacks. The core concept of triplet margin loss involves the use of triplets, each consisting of an anchor sample, a positive sample, and a negative sample. The anchor and positive samples represent similar instances, while the negative sample represents a dissimilar one. The algorithm learns to embed these samples such that the distance between the anchor and positive sample is smaller than the distance between the anchor and negative sample.
Given training samples $x_{a}$, $x_{p}$, and $x_{n}$, they represent anchor, positive, and negative points, respectively. The RevPRAG embedding model will output closer embedding vectors for any $Act_{a}$ and $Act_{p}$, but farther for any $Act_{a}$ and $Act_{n}$. The loss function is formally defined as: $L=\max(\text{Dist}\left(Act_{a},Act_{p}\right)-\text{Dist}\left(Act_{a},Act_{n}\right)+margin,0)$, where $\text{Dist}(\cdot,\cdot)$ denotes a distance metric (typically the Euclidean distance), and margin is a positive constant. The presence of the margin introduces an additional parameter in triplet loss that requires tuning. If the margin is too large, the model’s loss will be high, and it may struggle to converge to zero. However, a larger margin generally improves the model’s ability to distinguish between very similar samples, making it easier to differentiate between $Act_{p}$ and $Act_{n}$. Conversely, if the margin is too small, the loss will quickly approach zero, making the model easier to train, but it will be less effective at distinguishing between $Act_{p}$ and $Act_{n}$.
At test time, given a test sample $x_{t}$, we compute the distance between its embedding $Act_{x_{t}}$ and the embedding of the support sample $Act_{x_{s}},x_{s}\in S$. The support set $S$ refers to a constructed dataset comprising labeled data that is excluded from the training set, denoted as $\left\{x_{s_{1}},...,x_{s_{n}}\right\}$, and corresponding labels are $\left\{T_{x_{s_{1}}},...,T_{x_{s_{n}}}\right\}$. It provides a reference for comparison and classification of new, unseen test data. The main purpose of the support set is to help determine labels for the test data. The label of the test data $x_{t}$ will be determined according to the label of the support sample $x_{s}$ that is closest to it. That is, $x_{t}$ is assigned the label of $x_{s}$, meaning $T_{x_{t}}=T_{x_{s}},$ where $x_{s}=argmin_{i}Dist(Act_{x_{t}},Act_{x_{s_{i}}})$. Here, $x_{s}$ is the nearest support data to the test data $x_{t}$.
Table 1: RevPRAG achieved high TPRs and low FPRs on three datasets for RAG with five different LLMs.
Dataset Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B Llama2-13B
NQ TPR 0.982 0.994 0.985 0.986 0.989
FPR 0.006 0.006 0.019 0.009 0.019
HotpotQA TPR 0.972 0.985 0.977 0.973 0.970
FPR 0.016 0.061 0.022 0.017 0.070
MS-MARCO TPR 0.988 0.989 0.999 0.978 0.993
FPR 0.007 0.012 0.001 0.011 0.025
5 Evaluation
5.1 Experimental Setup
RAG Setup. RAG comprises three key components: knowledge database, retriever, and LLM. The setup is shown below:
$\bullet$ Knowledge Database: We leverage three representative benchmark question-answering datasets in our evaluation: Natural Questions (NQ) Kwiatkowski et al. (2019), HotpotQA Yang et al. (2018), MS-MARCO Bajaj et al. (2016). Please note that RevPRAG can be expanded to cover poisoning attacks towards any other datasets used for RAG systems, not limited to the datasets used in this paper. To evaluate the detection of poisoning attacks to the knowledge database of RAG, we selected 3,000 instances of the triple (q, t, a) from each of the above three evaluation datasets. Among them, 70% were used for training, 20% for testing, and 10% as the support dataset. In each triple, q is the question, t is the texts collected from Wikipedia or web documents corresponding to q, and a is the correct answer to the question q, generated using a state-of-the-art LLM.
To better simulate the RAG poisoning attack scenario implemented through the knowledge database, we will employ three different methods for generating poisoning texts in the experiments, including PoisonedRAG Zou et al. (2024) , GARAG Cho et al. (2024), and PRCAP Zhong et al. (2023). We will evaluate the performance of our proposed detection approach across this series of different scenarios.
$\bullet$ Retriever: In our experiments, we evaluate four state-of-the-art dense retrieval models: Contriever Izacard et al. (2021) (pre-trained), Contriever-ms (fine-tuned on MS-MARCO) Izacard et al. (2021), DPR-mul Karpukhin et al. (2020) (trained on multiple datasets), and ANCE Xiong et al. (2020) (trained on MS-MARCO). Based on previous works Lewis et al. (2020),  Zhong et al. (2023), we default to using the dot product between the embedding vectors of a question and a document in the knowledge database to calculate their similarity score, and we retrieve the five most similar texts from the knowledge database to serve as the context for a given question.
$\bullet$ LLM: Our experiments are conducted on several popular LLMs, each with distinct architectures and characteristics. We employ GPT2-XL 1.5B Radford et al. (2019), Llama2-7B Touvron et al. (2023), Llama2-13B, Mistral-7B Jiang et al. (2023), Llama3-8B, and Llama2-13B. These LLMs were chosen for their open-source nature and to facilitate comparisons with other methodologies. We use the prompt shown in Fig. 5 to guide the LLMs in generating an answer to a question.
Baselines. To the best of our knowledge, there are currently no dedicated detection methods or evaluations specifically for RAG poisoning attacks. Thus, we extend two current methods  Li et al. (2024) and  Xi et al. (2024) for the RAG scenario. CoS Li et al. (2024) is a black-box approach that guides the LLM to generate detailed reasoning steps for the input, subsequently scrutinizing the reasoning process to ensure consistency with the final answer. MDP Xi et al. (2024) is a white-box method that exploits the disparity in masking sensitivity between poisoned and clean samples.
Evaluation Metrics. The effectiveness of the proposed detection method is assessed using two metrics. The True Positive Rate (TPR), which measures the proportion of effectively poisoned responses that are successfully detected. The False Positive Rate (FPR), which quantifies the proportion of correct responses that are misclassified as being caused by poisoning attacks. Our primary goal is to detect poisoned responses as effectively as possible while minimizing the impact on RAG’s normal functionality, which is why we have selected these two metrics.
5.2 Overall Results
RevPRAG achieves high TPRs and low FPRs. Table 1 shows the TPRs and FPRs of RevPRAG on three datasets. We have the following observations from the experimental results. First, RevPRAG achieved high TPRs consistently on different datasets and LLMs when injecting five poisoned texts into the knowledge database. For instance, RevPRAG achieved 98.5% (on NQ), 97.7% (on HotpotQA), and 99.9% (on MS-MARCO) TPRs for RAG with Mistral-7B. Our experimental results show that assessing whether the output of a RAG system is correct or poisoned based on the activations of LLMs is both highly feasible and reliable (i.e., capable of achieving exceptional accuracy). Second, RevPRAG achieves low FPRs under different settings, e.g., close to 1% in nearly all cases. This result indicates that our approach not only maximizes the detection of poisoned responses but also maintains a low false positive rate, significantly reducing the risk of misclassifying correct answers as poisoned.
We also conduct experiments on different retrievers. Table 2 shows the TPRs and FPRs of RevPRAG on HotpotQA for RAG with different retrievers and LLMs. Results show that our approach consistently achieved high TPRs and low FPRs across RAG with various retrievers and LLMs. For instance, RevPRAG achieves 97.2% (with Contriever), 98.7% (with Contriever-ms), 97.9% (with DPR-mul), 97.8% (with ANCE) TPRs alongside 1.6% (with Contriever), 5.7% (with Contriever-ms), 3.5% (with DPR-mul), and 4.2% (with ANCE) FPRs for RAG when using GPT2-XL 1.5B.
Table 2: RevPRAG achieved high TPRs and low FPRs on HotpotQA for RAG with four different retrievers.
Attack Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B
Contriever TPR 0.972 0.985 0.977
FPR 0.016 0.061 0.022
Contriever-ms TPR 0.987 0.983 0.998
FPR 0.057 0.018 0.012
DPR-mul TPR 0.979 0.966 0.999
FPR 0.035 0.075 0.001
ANCE TPR 0.978 0.981 0.993
FPR 0.042 0.028 0.023
Table 3: RevPRAG outperforms baselines.
Dataset Metrics Methods
CoS MDP Ours
NQ TPR 0.488 0.946 0.986
FPR 0.146 0.108 0.009
HotpotQA TPR 0.194 0.886 0.973
FPR 0.250 0.372 0.017
MS-MARCO TPR 0.771 0.986 0.978
FPR 0.027 0.181 0.011
RevPRAG outperforms baselines. Table 3 compares RevPRAG with baselines for RAG with Llama3-8B under the default settings. We have the following observations. First, the Chain-of-Scrutiny (CoS), a backdoor detection method based on reasoning chain analysis, has demonstrated limited effectiveness. We attribute this to the fact that CoS is specifically designed for detecting backdoor attacks in LLMs, relying on the shortcut from the trigger to the target output. This differs from RAG, where backdoor attacks are carried out by injecting poisoned texts into the knowledge database. Second, MDP achieves good TPRs, but it also exhibits relatively high FPRs, reaching as much as 37.2%. However, MDP is an input-based approach that focuses on detecting whether the input is poisoned. In contrast, our approach concentrates on determining whether the responses generated by RAG are correct or poisoned, as we observe that the correctness (or not) of RAG’s responses provides greater robustness in indicating poisoning attacks.
Table 4: The TPRs and FPRs of RevPRAG for different poisoned text generation methods on HotpotQA.
Attack Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B
PoisonedRAG TPR 0.972 0.985 0.977
FPR 0.016 0.061 0.022
GARAG TPR 0.961 0.976 0.974
FPR 0.025 0.046 0.026
PRCAP TPR 0.966 0.986 0.965
FPR 0.012 0.061 0.022
5.3 Ablation Study
Different methods for generating poisoned texts. To ensure the effectiveness of the evaluation, we employ three different methods introduced by PoisonedRAG, GARAG, and PRCAP to generate the poisoned texts. The experimental results in Table 4 show that RevPRAG consistently achieves high TPRs and low FPRs when confronted with poisoned texts generated by different strategies. For instance, RevPRAG achieved 97.2% (with GPT2-XL 1.5B), 98.5% (with Llama2-7B), and 97.7% (with Mistral-7B) TPRs for poisoned texts generated with PoisonedRAG.
Table 5: The TPRs and FPRs of RevPRAG for different quantities of injected poisoned text on HotpotQA (total retrieved texts: five).
Quantity Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B
five TPR 0.972 0.985 0.977
FPR 0.016 0.061 0.022
four TPR 0.976 0.977 0.986
FPR 0.034 0.047 0.033
three TPR 0.963 0.986 0.995
FPR 0.011 0.043 0.004
two TPR 0.971 0.995 0.991
FPR 0.011 0.047 0.005
one TPR 0.970 0.988 0.989
FPR 0.049 0.031 0.022
Quantity of injected poisoned texts. Table 5 illustrates the impact of varying quantities of poisoned text on the detection performance of RevPRAG. The more poisoned texts are injected, the higher the likelihood of retrieving them for RAG processing. From the experimental results, we observe that even with varying amounts of injected poisoned text, RevPRAG consistently achieves high TPRs and low FPRs. For example, when the total number of retrieved texts is five and the injected quantity is two, RevPRAG achieves a 99.5% TPR and a 4.7% FPR for RAG with Llama2-7B. The reason for this phenomenon is that the similarity between the retrieved poisoned texts and the query is higher than that of clean texts. Consequently, the LLM generates responses based on the content of the poisoned texts.
Effects of different support set size. In RevPRAG, the Support Data plays a pivotal role in the model’s learning process. It supplies labeled information and task-specific knowledge, allowing the model to conduct effective reasoning and learning even with limited data. We experiment with various support set sizes ranging from 50 to 250 to examine their effect on the performance of RevPRAG. This evaluation was conducted on Llama2-7B with different datasets. The results in Fig. 6 indicate that varying the support size does not significantly impact the model’s detection performance.
Figure 6: Effects of support set size.
6 Conclusion
In this work, we find that correct and poisoned responses in RAG exhibit distinct differences in LLM activations. Building on this insight, we develop RevPRAG, a detection pipeline that leverages these activations to identify poisoned responses in RAG caused by the injection of malicious texts into the knowledge database. Our approach demonstrates robust performance across RAGs utilizing five different LLMs and four distinct retrievers on three datasets. Experimental results show that RevPRAG achieves exceptional accuracy, with true positive rates approaching 98% and false positive rates near 1%. Ablation studies further validate its effectiveness in detecting poisoned responses across different types and levels of poisoning attacks. Overall, our approach can accurately distinguish between correct and poisoned responses.
Limitations.
Our work has the following limitations:
  • This work does not propose a specific method for defending against poisoning attacks on RAG. Instead, our focus is on the timely detection of poisoned responses generated by the LLM, aiming to prevent potential harm to users from such attacks.
  • Our approach requires accessing the activations of the LLM, which necessitates the LLM being a white-box model. While this may present certain limitations for users, our method can be widely adopted by LLM service providers. Providers can implement our strategy to ensure the reliability of their services and enhance trust with their users.
  • Our approach primarily focuses on determining whether the response generated by the RAG is correct or poisoned, without delving into more granular distinctions. The main goal of our study is to protect users from the impact of RAG poisoning attacks, while more detailed classifications of RAG responses will be addressed in future work.
Ethics Statement
The goal of this work is to detect whether a RAG has generated a poisoned response. All the data used in this study is publicly available, so it does not introduce additional privacy concerns. All source code and software will be made open-source. While the open-source nature of the code may lead to adaptive attacks, we can further enhance our model by incorporating more internal and external information. Overall, we believe our approach can further promote the secure application of RAG.
Appendix A Additional Results
Table 6: Generalization Performance of RevPRAG for RAG with four different LLMs. The training and test datasets vary across different rows. Abbreviations: Hot (HotpotQA), MS (MS-MARCO).
Training Dataset Test Dataset Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
NQ & Hot MS TPR 0.881 0.886 0.948 0.956
FPR 0.134 0.149 0.076 0.066
Hot & MS NQ TPR 0.980 0.983 0.988 0.980
FPR 0.007 0.074 0.078 0.038
NQ & MS Hot TPR 0.977 0.961 0.942 0.978
FPR 0.025 0.089 0.055 0.049
NQ & Hot & MS NQ & Hot & MS TPR 0.986 0.994 0.985 0.987
FPR 0.032 0.007 0.009 0.035
A.1 Generalization
Given the wide range of RAG application scenarios and the diverse user requirements it faces, it is impractical to ensure that our detection model has been trained on all possible scenarios and queries in real-world applications. However, the performance of neural network models largely depends on the similarity between the distributions of the training data and the test data Yang et al. (2024). Consequently, our model’s performance may degrade when faced with training and test data that stem from differing distributions—a challenge frequently observed in real-world scenarios. To address this issue, we conduct a series of generalization experiments. Specifically, we train the detection model using any two datasets and test it on a third dataset that was not used during training. For example, we use NQ and HotpotQA as training datasets and MS-MARCO as the testing dataset. Although these three datasets are all QA datasets, they exhibit significant differences. For example, NQ focuses on extracting answers to factual questions from a single long document, HotpotQA involves multi-document reasoning to derive answers, and MS-MARCO retrieves and ranks relevant answers from a large-scale collection of documents. Therefore, conducting generalization experiments based on these three datasets is reasonable.
Table 6 illustrates the TPRs and FPRs of RevPRAG for RAG with four different LLMs. Overall, the experimental results demonstrate that our detection model exhibits strong generalization performance across RAG with different LLMs and various datasets. For example, when using HotpotQA and MS-MARCO as training data, the detection model achieves TPRs of 98% (with GPT2-XL 1.5B), 98.3% (with Llama2-7B), 98.8% (with Mistral-7B), and 98% (with Llama3-8B) on the NQ dataset. Meanwhile, all FPRs remain below 8%. Furthermore, we observe that the generalization performance is best when NQ is used as the test data (for instance, 98.3% with Llama2-7B), while MS-MARCO shows the poorest performance (for instance, 88.6% with Llama2-7B). We attribute this to the fact that the questions and tasks in HotpotQA and MS-MARCO are more complex compared to those in NQ. Therefore, detection models trained on more complex tasks generalize well to simpler tasks, whereas the reverse is more challenging. In conclusion, these experimental results highlight that RevPRAG exhibits strong generalization and robust detection performance, even in the presence of significant discrepancies between the training and test datasets.
(a) TPRs of RevPRAG.
(b) FPRs of RevPRAG.
Figure 7: RevPRAG trained on the activations from specific layers.
A.2 RevPRAG’s Performance on Complex Open-Ended Questions
In this section, we conducted a series of experiments to evaluate the performance of RevPRAG on complex, open-ended questions (e.g., “how to make relationship last?”). These questions present unique challenges due to their diverse and unstructured nature, in contrast to straightforward, closed-ended questions (e.g., “What is the name of the highest mountain?”). In our experiments, the NQ, HotpotQA, and MS-MARCO datasets primarily consist of close-ended questions. As a result, the majority of our previous experiments focused on close-ended problems, which was our default experimental setting. In this study, we utilized the advanced GPT-4o to filter and extract 3,000 open-ended questions from the HotpotQA and MS-MARCO datasets for training and testing the model. For open-ended questions, cosine similarity is employed to evaluate whether the LLM’s response aligns with the attacker’s target response. If the similarity surpasses a predefined threshold, it is considered indicative of a successful poisoning attack.
The experimental results are shown in Table 7. We can observe that RevPRAG demonstrates excellent detection performance even on complex open-ended questions. For example, RevPRAG achieved TPRs of 99.1% on HotpotQA and 99.0% on MS-MARCO, alongside FPRs of 0.8% on HotpotQA and 0.1% on MS-MARCO for RAG utilizing the Mistral-7B model.
Table 7: RevPRAG achieved high TPRs and low FPRs on the open-ended questions from HotpotQA and MS-MARCO datasets.
Dataset Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
HotpotQA TPR 0.982 0.995 0.991 0.982
FPR 0.033 0.029 0.008 0.007
MS-MARCO TPR 0.988 0.989 0.990 0.983
FPR 0.009 0.009 0.001 0.017
Table 8: RevPRAG trained on the activations from specific individual layers of GPT2-XL 1.5B.
Dataset Metrics Different layers
layer 0 layer 8 layer 15 layer 22 layer 29 layer 36 layer 41 layer 47
NQ TPR 0.996 0.988 0.996 0.984 0.996 0.988 0.992 0.996
FPR 0.027 0.007 0.017 0.003 0.007 0.003 0.017 0.003
HotpotQA TPR 0.713 0.984 0.994 0.989 0.520 0.619 0.931 0.992
FPR 0.409 0.023 0.012 0.006 0.445 0.409 0.023 0.019
MS-MARCO TPR 0.967 0.998 0.988 0.986 0.988 0.963 0.955 0.992
FPR 0.023 0.004 0.002 0.019 0.030 0.026 0.037 0.017
A.3 Activations from Specified Layers.
Fig. 7 illustrates the detection performance of RevPRAG using activations from different layers of various LLMs. In previously presented experiments, we utilize activations from all layers as both training and testing data, yielding excellent results. Additionally, we also test using different layers. The experimental results in Fig. 7 demonstrate that utilizing activations from only the first few layers can still achieve satisfactory detection performance, providing valuable insights for future research. For example, when using activations from layers 0 to 5, RevPRAG achieved TPRs exceeding 97% while maintaining FPRs below 7% for RAG with all LLMs on HotpotQA. However, the experimental results also suggest that using activations from intermediate or deeper layers can lead to performance fluctuations, including signs of degradation or slower convergence. For instance, when using activations from layers 16 to 24 with Llama3-8B as the LLM in RAG, RevPRAG achieves a TPR of 78.8% on NQ dataset and 86% on MS-MARCO dataset.
We further explored the use of activations from a specific individual layer of the LLMs to train and test RevPRAG. We chose 8 layers with roughly even spacing for testing. As shown in Table 8, when using activations from only a specific layer of the GPT2-XL model, RevPRAG demonstrates excellent performance in general. For instance, when the model is trained using activations from layer 0 on the NQ dataset, the TPR can reach as high as 99.6%. However, we also observed that activations from certain layers do not yield satisfactory performance. For example, when the model is trained using activations from layer 29 on the HotpotQA dataset, the TPR is only 52%, while the FPR reaches 44.5%. It is precisely due to the existence of these suboptimal layers that models trained with multi-layer activations may not always outperform those using single-layer activations (such as layer 0 with NQ dataset). However, incorporating multi-layer activations can enhance the model’s stability, mitigating the detrimental effects of these suboptimal layers.
A.4 Impact of Similarity Metric.
Table 9 shows the results on the HotpotQA dataset, indicating that the choice of similarity calculation method has minimal impact on RevPRAG’s performance, which consistently achieves high TPR and low FPR. This suggests the robustness of our approach, as it reliably identifies poisoned texts even when LLM activations vary slightly under similar conditions.
Table 9: Impact of similarity metric.
Similarity Metric Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
Dot Product TPR 0.972 0.985 0.977 0.973
FPR 0.016 0.061 0.022 0.017
Cosine TPR 0.978 0.990 0.979 0.981
FPR 0.037 0.011 0.023 0.043
A.5 Isolating Poisoned Responses and Hallucinations
It is well-known that hallucinations are an inevitable phenomenon in LLMs. Even with the introduction of a knowledge database in RAG, LLMs may still generate non-factual responses due to hallucinations. Therefore, the incorrect responses generated by RAG may also stem from hallucinations, rather than being solely caused by RAG poisoning. We conducted experiments to test if our approach can distinguish hallucinations and RAG poisoning. Fig. 8 shows the t-SNE representation of mean activations for poisoned response and hallucinations across all layers for Mistral-7B and Llama2-7B on the NQ dataset. We observe that activations across all layers clearly distinguish between hallucinations and poisoned responses.
Figure 8: t-SNE visualizations of activations for poisoned responses and hallucinations.
This key finding has led us to extend our approach to differentiate between poisoned responses and hallucinations. We thus continue to collect data and train the model using the process outlined in Fig. 3, with the only difference being that we now collect hallucination data. We also conduct extensive experiments on RAG with different LLMs and datasets. From the experimental results in Table 10, we can see that our method achieves a high TPR across all LLMs and datasets. For instance, RevPRAG achieved 98.7% (on NQ), 97.5% (on HotpotQA), and 97.3% (on MS-MARCO) TPRs for RAG with GPT2-XL 1.5B. Furthermore, we observe that the FPR remains low across all evaluation settings. As shown in the table, RevPRAG could achieve 0.8% (on NQ), 0.8% (on HotpotQA) and 0.6% (on MS-MARCO) FPRs for RAG with Llama3-8B. This further supports our previous observation that there is a clear distinction between poisoned responses and hallucinations.
Table 10: RevPRAG could achieve high TPRs and low FPRs to distinguish poisoned responses and hallucinations.
Dataset Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
NQ TPR 0.987 0.983 0.993 0.995
FPR 0.046 0.017 0.069 0.008
HotpotQA TPR 0.975 0.978 0.991 0.995
FPR 0.004 0.058 0.004 0.008
MS-MARCO TPR 0.973 0.984 0.999 0.989
FPR 0.009 0.023 0.001 0.006

### v4

@@
 RevPRAG: Revealing Poisoning Attacks in Retrieval-Augmented Generation through LLM Activation Analysis
-Xue Tan1, Hao Luan1, Mingyu Luo1, Xiaoyan Sun2, Ping Chen3, Jun Dai2 1School of Computer Science, Fudan University, Shanghai, China, 2Department of Computer Science, Worcester Polytechnic Institute, MA, USA, 3Institute of Big Data, Fudan University, Shanghai, China, Correspondence: xsun7@wpi.edu pchen@fudan.edu.cn jdai@wpi.edu
+Xue Tan1,3, Hao Luan1,3, Mingyu Luo1,3, Xiaoyan Sun2, Ping Chen3, Jun Dai2 1School of Computer Science, Fudan University, Shanghai, China 2Department of Computer Science, Worcester Polytechnic Institute, MA, USA 3Institute of Big Data, Fudan University, Shanghai, China Correspondence: pchen@fudan.edu.cn, xsun7@wpi.edu, jdai@wpi.edu
 1 Introduction
@@
 (2024).
-RAG poisoning refers to the act of injecting malicious or misleading content into the knowledge database, contaminating the retrieved texts in RAG and ultimately leading it to produce the attacker’s desired response (e.g., the target answer could be “Fuji" when the target question is “What is the name of the highest mountain?").
+RAG poisoning refers to the act of injecting malicious or misleading content into the knowledge database, contaminating the retrieved texts and ultimately leading the LLM to produce the attacker’s desired response (e.g., the target answer could be “Fuji" when the target question is “What is the name of the highest mountain?").
 This attack leverages the dependency between LLMs and the knowledge database, transforming the database into a new attack surface to facilitate poisoning.
@@
 Specifically, we empirically analyze the activations of the final token in the input sequence across all layers of the LLM.
-Our findings demonstrate that it is highly feasible to differentiate between these two types of responses by comparing the activations of the LLM when generating correct responses versus poisoned ones.
+Our findings demonstrate that the model exhibits distinguishable activation patterns when generating correct versus poisoned responses.
 Based on this, we propose a systematic and automated detection pipeline, namely RevPRAG, which consists of three key components: poisoned data collection, LLM activation collection and preprocessing, and the detection model design.
@@
 (2020).
-In this work, to defend against RAG attacks, we present a detection mechanism that can promptly capture key information during the model’s response generation process and assess whether the current response is trustworthy or not.
+LLM Factoscope He et al.
+(2024) is a runtime detection tool that leverages the internal states of LLMs, such as activation maps, output rankings, and top-$k$ probabilities, to identify factual inaccuracies caused by model hallucinations.
+While Factoscope is effective at detecting hallucinations in general LLMs, it is not designed to address RAG poisoning attacks, which result from manipulations of the external knowledge base rather than internal model errors.
+Its complex architecture with multiple sub-models makes it less suitable for latency-sensitive RAG applications.
+In this work, we present RevPRAG, a method that addresses these gaps by: (1) focusing on RAG-specific poisoning attacks and conducting extensive tests to validate its effectiveness in detecting such attacks (Section 5), (2) using a lightweight, activation-based pipeline optimized for real-time detection of whether an RAG response is trustworthy (Section 5.6), and (3) evaluations show that our performance (Section 5.2) and efficiency (Section 5.6) surpass those of Factoscope.
 3 Preliminary
@@
 Figure 2: t-SNE visualizations of activations for correct and poisoned responses.
-Figure 3: The workflow of RevPRAG.
 3.2 Rationale
@@
 This clearly demonstrates the distinguishability between the two types of responses, to some extent, supports our conjecture.
-Figure 4: An instance of using RevPRAG.
+Figure 3: The workflow of RevPRAG.
 4 Methodology
@@
 (2023)) for generating poisoned texts and increasing the similarity between the poisoned texts and the queries, to raise the likelihood that the poisoned texts would be selected by the retriever.
-The retrieved texts, along with the question, are then used to construct a new prompt for the LLM to generate the answer.
-The prompt Zou et al.
-(2024), as shown in Fig.
-5, is employed to achieve this objective.
-Figure 5: The prompt used in RAG to make an LLM generate an answer based on the retrieved texts.
+A detailed introduction of these methods can be found in Section A.2.
+The retrieved texts and the question are combined into a new prompt, following the format in Zou et al.
+(2024) (see Fig.
+6 in Section A.3), for LLM answer generation.
+Figure 4: An instance of using RevPRAG.
 4.3 Activation Collection and Processing
-For an LLM input sequence $X=\left(x_{1},x_{2},\cdots,x_{n}\right)$, we extract the activations $Act_{n}$ for the last token $x_{n}$ in the input across all layers in the LLM as a summary of the context.
+For an LLM input sequence $X=\left(t_{1},t_{2},\cdots,t_{n}\right)$, we extract the activations $Act_{n}$ for the last token $x_{n}$ in the input across all layers in the LLM as a summary of the context.
 The activations $Act_{n}$ contain the inner representations of the LLM’s knowledge related to the input.
@@
 (2023).
-As mentioned earlier, there is a significant difference between the LLM activations for the poisoned responses and the activations for correct responses.
+We collect two types of activations: correct activations (labeled as 1), obtained when the LLM retrieves accurate content and generates the correct response; and poisoned activations (labeled as 0), obtained when the LLM retrieves poisoned content and produces the attacker’s target response.
 We introduce normalization of the activations for effective integration into the training process.
-We calculate the mean $\mu$ and standard deviation $\sigma$ of the dataset.
-Then, we use the obtained $\mu$ and $\sigma$ to normalize the activations with the formula: $Act^{nor}_{n}=\left(Act_{n}-\mu\right)/\sigma$.
-This standardization improves training efficiency by scaling activations uniformly, preventing bias towards larger features.
-It also ensures smoother optimization, alleviates gradient-related issues, and enhances the performance of algorithms that rely on distance metrics.
+We calculate the mean $\mu$ and standard deviation $\sigma$ of the activations across all instances in the dataset.
+Then, we use the obtained $\mu$ and $\sigma$ to normalize the activations with the formula:
+$Act^{nor}_{n}=\left(Act_{n}-\mu\right)/\sigma.$ (1)
 4.4 RevPRAG Model Design
-After collecting and preprocessing the dataset of activations, we design a probing mechanism based on the dataset.
-Inspired by few-shot learning and Siamese networks, our proposed RevPRAG model is adept at distinguishing correct answers from poisoned ones and demonstrates strong generalizability from limited data.
-In order to efficiently capture the relationships between and within different layers of the LLM, we utilize Convolutional Neural Networks (CNNs) with the ResNet18 architecture He et al.
+After collecting and preprocessing the activation dataset, we partition it into a training set $D_{train}$, a test set $D_{test}$, and a support set $S$ to facilitate the construction and evaluation of the probe model.
+Drawing inspiration from few-shot learning and Siamese networks, the proposed RevPRAG model is designed to effectively distinguish between clean and poisoned responses, while demonstrating strong generalization capabilities even under limited data conditions.
+To efficiently capture both intra-layer and inter-layer relationships within the LLM, we employ Convolutional Neural Networks (CNNs) based on the ResNet18 architecture He et al.
 (2016).
-We use triplet networks which share the same architecture and weights to learn embeddings of the tasks as shown in Fig.
+Additionally, we adopt a triplet network structure, in which three subnetworks with shared architecture and weights are used to learn task embeddings, as illustrated in Fig.
 3.
@@
 (2015), a commonly used approach for tasks where it is difficult to distinguish similar instances.
-Triplet margin loss is a loss function used in training neural networks for tasks such as face recognition or object classification.
-At the same time, triplet margin loss is also widely used in few-shot learning scenarios.
-When model encounters out-of-distribution data, it can quickly adapt with a small amount of support data, without the need to retrain the entire model.
-The goal of this loss function is to learn a similarity metric within a high-dimensional embedding space (also known as feature space), where representations of similar objects (e.g., images of the same person) are close to each other, while representations of dissimilar objects are farther apart.
-This powerful similarity metric provided by triplet margin loss is particularly suitable for distinguishing LLM activations, enabling RevPRAG to effectively differentiate activation differences caused by poisoning attacks.
-The core concept of triplet margin loss involves the use of triplets, each consisting of an anchor sample, a positive sample, and a negative sample.
-The anchor and positive samples represent similar instances, while the negative sample represents a dissimilar one.
-The algorithm learns to embed these samples such that the distance between the anchor and positive sample is smaller than the distance between the anchor and negative sample.
-Given training samples $x_{a}$, $x_{p}$, and $x_{n}$, they represent anchor, positive, and negative points, respectively.
-The RevPRAG embedding model will output closer embedding vectors for any $Act_{a}$ and $Act_{p}$, but farther for any $Act_{a}$ and $Act_{n}$.
-The loss function is formally defined as: $L=\max(\text{Dist}\left(Act_{a},Act_{p}\right)-\text{Dist}\left(Act_{a},Act_{n}\right)+margin,0)$, where $\text{Dist}(\cdot,\cdot)$ denotes a distance metric (typically the Euclidean distance), and margin is a positive constant.
-The presence of the margin introduces an additional parameter in triplet loss that requires tuning.
-If the margin is too large, the model’s loss will be high, and it may struggle to converge to zero.
-However, a larger margin generally improves the model’s ability to distinguish between very similar samples, making it easier to differentiate between $Act_{p}$ and $Act_{n}$.
-Conversely, if the margin is too small, the loss will quickly approach zero, making the model easier to train, but it will be less effective at distinguishing between $Act_{p}$ and $Act_{n}$.
-At test time, given a test sample $x_{t}$, we compute the distance between its embedding $Act_{x_{t}}$ and the embedding of the support sample $Act_{x_{s}},x_{s}\in S$.
-The support set $S$ refers to a constructed dataset comprising labeled data that is excluded from the training set, denoted as $\left\{x_{s_{1}},...,x_{s_{n}}\right\}$, and corresponding labels are $\left\{T_{x_{s_{1}}},...,T_{x_{s_{n}}}\right\}$.
+The training data is randomly divided into triplets consisting of an anchor instance $x_{a}$, a positive instance $x_{p}$, and a negative instance $x_{n}$, where the anchor and positive belong to the same class, while the anchor and negative come from different classes.
+The triplet margin loss function is formally defined as:
+$L=\max\big{(}$ $\text{Dist}(x_{a},x_{p})-\text{Dist}(x_{a},x_{n})$
+$+\textit{\text{margin}},\ 0\big{)},$ (2)
+where $\text{Dist}(\cdot,\cdot)$ denotes a distance metric (typically the Euclidean distance), and margin is a positive constant.
+The training objective is to encourage the RevPRAG embedding model to output closer embedding vectors for any $x_{a}$ and $x_{p}$, but farther for any $x_{a}$ and $x_{n}$.
+At test time, given a test sample $x_{t}$, we compute the distance between its embedding and the embedding of the support sample $x_{s},x_{s}\in S$.
+The support set $S$ refers to a dataset comprising labeled data, denoted as $\left\{x_{s_{1}},...,x_{s_{n}}\right\}$, and corresponding labels are $\left\{T_{x_{s_{1}}},...,T_{x_{s_{n}}}\right\}$.
 It provides a reference for comparison and classification of new, unseen test data.
@@
 The label of the test data $x_{t}$ will be determined according to the label of the support sample $x_{s}$ that is closest to it.
-That is, $x_{t}$ is assigned the label of $x_{s}$, meaning $T_{x_{t}}=T_{x_{s}},$ where $x_{s}=argmin_{i}Dist(Act_{x_{t}},Act_{x_{s_{i}}})$.
+That is, $x_{t}$ is assigned the label of $x_{s}$, meaning $T_{x_{t}}=T_{x_{s}},$ where $x_{s}=argmin_{i}Dist(x_{t},x_{s_{i}})$.
 Here, $x_{s}$ is the nearest support data to the test data $x_{t}$.
@@
 Please note that RevPRAG can be expanded to cover poisoning attacks towards any other datasets used for RAG systems, not limited to the datasets used in this paper.
-To evaluate the detection of poisoning attacks to the knowledge database of RAG, we selected 3,000 instances of the triple (q, t, a) from each of the above three evaluation datasets.
-Among them, 70% were used for training, 20% for testing, and 10% as the support dataset.
-In each triple, q is the question, t is the texts collected from Wikipedia or web documents corresponding to q, and a is the correct answer to the question q, generated using a state-of-the-art LLM.
-To better simulate the RAG poisoning attack scenario implemented through the knowledge database, we will employ three different methods for generating poisoning texts in the experiments, including PoisonedRAG Zou et al.
-(2024) , GARAG Cho et al.
-(2024), and PRCAP Zhong et al.
-(2023).
-We will evaluate the performance of our proposed detection approach across this series of different scenarios.
+The detailed usage instructions for the dataset are provided in Section A.1.
 $\bullet$ Retriever: In our experiments, we evaluate four state-of-the-art dense retrieval models: Contriever Izacard et al.
@@
 (2020) (trained on MS-MARCO).
-Based on previous works Lewis et al.
-(2020),  Zhong et al.
-(2023), we default to using the dot product between the embedding vectors of a question and a document in the knowledge database to calculate their similarity score, and we retrieve the five most similar texts from the knowledge database to serve as the context for a given question.
-$\bullet$ LLM: Our experiments are conducted on several popular LLMs, each with distinct architectures and characteristics.
-We employ GPT2-XL 1.5B Radford et al.
+$\bullet$ LLM: Our experiments are conducted on several popular LLMs, each with distinct architectures and characteristics, including GPT2-XL 1.5B Radford et al.
 (2019), Llama2-7B Touvron et al.
 (2023), Llama2-13B, Mistral-7B Jiang et al.
-(2023), Llama3-8B, and Llama2-13B.
-These LLMs were chosen for their open-source nature and to facilitate comparisons with other methodologies.
-We use the prompt shown in Fig.
-5 to guide the LLMs in generating an answer to a question.
+(2023), and Llama3-8B.
+Unless otherwise specified, we adopt the following default settings: HotpotQA as the knowledge base, Contriever as the retriever, GPT2-XL 1.5B as the LLM, and 100 support samples.
+Moreover, we use the dot product between the embedding vectors of a question and a text to measure their similarity.
+Poisoned texts are generated following PoisonedRAG Zou et al.
+(2024).
+Consistent with prior work Lewis et al.
+(2020), we retrieve the 5 most similar texts from the knowledge database to serve as context for a given question.
 Baselines.
-To the best of our knowledge, there are currently no dedicated detection methods or evaluations specifically for RAG poisoning attacks.
-Thus, we extend two current methods  Li et al.
-(2024) and  Xi et al.
-(2024) for the RAG scenario.
+We compared RevPRAG with five existing methods, and although they were not specifically designed for detecting RAG poisoning attacks, we investigated their potential applications in this domain.
 CoS Li et al.
@@
 (2024) is a white-box method that exploits the disparity in masking sensitivity between poisoned and clean samples.
+LLM Factoscope He et al.
+(2024) leverages the internal states of LLMs to detect hallucinations, and we investigate its use for identifying poisoning attacks in RAG systems.
+Both RoBERTa Pan et al.
+(2023) and Discern Hong et al.
+(2024) employ an additional discriminator to distinguish whether the content retrieved by RAG consists of accurate documents or those that contradict factual information.
 Evaluation Metrics.
-The effectiveness of the proposed detection method is assessed using two metrics.
-The True Positive Rate (TPR), which measures the proportion of effectively poisoned responses that are successfully detected.
-The False Positive Rate (FPR), which quantifies the proportion of correct responses that are misclassified as being caused by poisoning attacks.
-Our primary goal is to detect poisoned responses as effectively as possible while minimizing the impact on RAG’s normal functionality, which is why we have selected these two metrics.
+$\bullet$ The True Positive Rate (TPR), which measures the proportion of effectively poisoned responses that are successfully detected.
+A higher TPR signifies better detection performance for poisoned responses, with a correspondingly lower rate of missed detections (i.e., lower false negative rate).
+$\bullet$ The False Positive Rate (FPR), which quantifies the proportion of correct responses that are misclassified as being caused by poisoning attacks.
+A lower FPR indicates fewer false positives for correct answers, minimizing disruption to the normal operation of RAG.
+Our goal is to detect poisoned responses as effectively as possible while minimizing the impact on RAG’s normal functionality, which is why we have selected these two metrics.
 5.2 Overall Results
@@
 This result indicates that our approach not only maximizes the detection of poisoned responses but also maintains a low false positive rate, significantly reducing the risk of misclassifying correct answers as poisoned.
-We also conduct experiments on different retrievers.
-Table 2 shows the TPRs and FPRs of RevPRAG on HotpotQA for RAG with different retrievers and LLMs.
-Results show that our approach consistently achieved high TPRs and low FPRs across RAG with various retrievers and LLMs.
-For instance, RevPRAG achieves 97.2% (with Contriever), 98.7% (with Contriever-ms), 97.9% (with DPR-mul), 97.8% (with ANCE) TPRs alongside 1.6% (with Contriever), 5.7% (with Contriever-ms), 3.5% (with DPR-mul), and 4.2% (with ANCE) FPRs for RAG when using GPT2-XL 1.5B.
 Table 2: RevPRAG achieved high TPRs and low FPRs on HotpotQA for RAG with four different retrievers.
@@
 Table 3: RevPRAG outperforms baselines.
-Dataset Metrics Methods
-CoS MDP Ours
-NQ TPR 0.488 0.946 0.986
-FPR 0.146 0.108 0.009
-HotpotQA TPR 0.194 0.886 0.973
-FPR 0.250 0.372 0.017
-MS-MARCO TPR 0.771 0.986 0.978
-FPR 0.027 0.181 0.011
+Dataset Metrics Baselines and Our Method
+CoS Li et al. (2024) MDP Xi et al. (2024) LLM Factoscope He et al. (2024) RoBERTa Pan et al. (2023) Discern Hong et al. (2024) Ours
+NQ TPR 0.488 0.946 0.949 0.977 0.810 0.986
+FPR 0.146 0.108 0.033 0.063 0.112 0.009
+HotpotQA TPR 0.194 0.886 0.939 0.956 0.817 0.973
+FPR 0.250 0.372 0.021 0.018 0.101 0.017
+MS-MARCO TPR 0.771 0.986 0.945 0.946 0.795 0.978
+FPR 0.027 0.181 0.028 0.070 0.101 0.011
+We also conduct experiments on different retrievers.
+Table 2 shows that our approach consistently achieved high TPRs and low FPRs across RAG with various retrievers and LLMs.
+For instance, RevPRAG achieves 97.2% (with Contriever), 98.7% (with Contriever-ms), 97.9% (with DPR-mul), 97.8% (with ANCE) TPRs alongside 1.6% (with Contriever), 5.7% (with Contriever-ms), 3.5% (with DPR-mul), and 4.2% (with ANCE) FPRs for RAG when using GPT2-XL 1.5B.
 RevPRAG outperforms baselines.
-Table 3 compares RevPRAG with baselines for RAG with Llama3-8B under the default settings.
-We have the following observations.
-First, the Chain-of-Scrutiny (CoS), a backdoor detection method based on reasoning chain analysis, has demonstrated limited effectiveness.
-We attribute this to the fact that CoS is specifically designed for detecting backdoor attacks in LLMs, relying on the shortcut from the trigger to the target output.
-This differs from RAG, where backdoor attacks are carried out by injecting poisoned texts into the knowledge database.
-Second, MDP achieves good TPRs, but it also exhibits relatively high FPRs, reaching as much as 37.2%.
-However, MDP is an input-based approach that focuses on detecting whether the input is poisoned.
-In contrast, our approach concentrates on determining whether the responses generated by RAG are correct or poisoned, as we observe that the correctness (or not) of RAG’s responses provides greater robustness in indicating poisoning attacks.
+Table 3 compares RevPRAG with baselines for RAG using Llama3-8B under the default settings.
+The overall results demonstrate the superiority of our approach.
+Meanwhile, several key observations can be drawn from the comparison.
+First, the limited effectiveness of CoS Li et al.
+(2024) may stem from its design focus on detecting backdoor attacks in LLMs via trigger-to-output shortcuts, which differs from RAG’s attack surface involving poisoned knowledge base entries.
+Second, MDP Xi et al.
+(2024) achieves good TPRs, but it also exhibits relatively high FPRs, reaching as much as 37.2%.
+LLM Factoscope He et al.
+(2024) leverages multiple internal states of LLMs, relying on layer-wise consistency for effective hallucination detection.
+However, it may not be suitable for targeted attacks like poisoning, and the use of diverse state data increases computational overhead and discriminator model complexity (Section 5.6).
+Input-based methods such as MDP Xi et al.
+(2024), RoBERTa Pan et al.
+(2023), and Discern Hong et al.
+(2024) aim to detect whether the input is poisoned.
+In contrast, our method focuses on determining whether the responses generated by RAG are correct or poisoned, as response correctness offers a more robust signal of poisoning attacks.
+5.3 Ablation Study
+Different methods for generating poisoned texts.
+To ensure the effectiveness of the evaluation, we employ three different methods introduced by PoisonedRAG, GARAG, and PRCAP to generate the poisoned texts.
+The experimental results in Table 4 show that RevPRAG consistently achieves high TPRs and low FPRs when confronted with poisoned texts generated by different strategies.
+For instance, RevPRAG achieved 97.2% (with GPT2-XL 1.5B), 98.5% (with Llama2-7B), and 97.7% (with Mistral-7B) TPRs for poisoned texts generated with PoisonedRAG.
 Table 4: The TPRs and FPRs of RevPRAG for different poisoned text generation methods on HotpotQA.
@@
 FPR 0.012 0.061 0.022
-5.3 Ablation Study
-Different methods for generating poisoned texts.
-To ensure the effectiveness of the evaluation, we employ three different methods introduced by PoisonedRAG, GARAG, and PRCAP to generate the poisoned texts.
-The experimental results in Table 4 show that RevPRAG consistently achieves high TPRs and low FPRs when confronted with poisoned texts generated by different strategies.
-For instance, RevPRAG achieved 97.2% (with GPT2-XL 1.5B), 98.5% (with Llama2-7B), and 97.7% (with Mistral-7B) TPRs for poisoned texts generated with PoisonedRAG.
 Table 5: The TPRs and FPRs of RevPRAG for different quantities of injected poisoned text on HotpotQA (total retrieved texts: five).
@@
 Effects of different support set size.
-In RevPRAG, the Support Data plays a pivotal role in the model’s learning process.
-It supplies labeled information and task-specific knowledge, allowing the model to conduct effective reasoning and learning even with limited data.
+In RevPRAG, support data provides essential labeled and task-specific information, facilitating effective reasoning and learning under limited data conditions.
 We experiment with various support set sizes ranging from 50 to 250 to examine their effect on the performance of RevPRAG.
-This evaluation was conducted on Llama2-7B with different datasets.
 The results in Fig.
@@
 Figure 5: Effects of support set size.
+5.4 RevPRAG’s Performance on Complex Open-Ended Questions
+In this section, we conducted a series of experiments to evaluate the performance of RevPRAG on complex, open-ended questions (e.g., “how to make relationship last?”).
+These questions present unique challenges due to their diverse and unstructured nature, in contrast to straightforward, closed-ended questions (e.g., “What is the name of the highest mountain?”).
+In our experiments, the NQ, HotpotQA, and MS-MARCO datasets primarily consist of close-ended questions.
+As a result, the majority of our previous experiments focused on close-ended problems, which was our default experimental setting.
+In this study, we utilized the advanced GPT-4o to filter and extract 3,000 open-ended questions from the HotpotQA and MS-MARCO datasets for training and testing the model.
+For open-ended questions, cosine similarity is employed to evaluate whether the LLM’s response aligns with the attacker’s target response.
+If the similarity surpasses a predefined threshold, it is considered indicative of a successful attack.
+Table 6: RevPRAG achieved high TPRs and low FPRs on the open-ended questions from HotpotQA and MS-MARCO datasets.
+Dataset Metrics LLMs of RAG
+GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
+HotpotQA TPR 0.982 0.995 0.991 0.982
+FPR 0.033 0.029 0.008 0.007
+MS-MARCO TPR 0.988 0.989 0.990 0.983
+FPR 0.009 0.009 0.001 0.017
+The experimental results are shown in Table 6.
+We can observe that RevPRAG demonstrates excellent detection performance even on complex open-ended questions.
+For example, RevPRAG achieved TPRs of 99.1% on HotpotQA and 99.0% on MS-MARCO, alongside FPRs of 0.8% on HotpotQA and 0.1% on MS-MARCO for RAG utilizing the Mistral-7B model.
+5.5 Effect of Real-world Natural Text Noise on Detection Performance
+To better approximate real-world scenarios, we injected natural textual noise (e.g., spelling and grammatical errors) into clean texts using the GPT-4o model.
+Noise was introduced in a controlled manner to preserve semantics while simulating realistic imperfections.
+Specifically, 50% of the clean texts were modified, with noise accounting for 10% of the total word count in each case.
+Notably, noise injection was applied only to clean texts, as poisoned texts are typically well-formed to maximize attack effectiveness and thus unlikely to contain such artifacts.
+As shown in Table 7, the experimental results indicate that injecting a certain amount of natural noise into clean texts does not significantly affect the detection performance of our proposed method.
+This further demonstrates the robustness of our approach.
+We attribute this to two main factors.
+First, the textual perturbations introduced by poisoning attacks are fundamentally different in nature from natural noise, and this distinction is clearly reflected in the activation vectors of the LLM.
+Second, LLMs are inherently robust to natural noise, which does not substantially interfere with their ability to generate correct responses.
+Table 7: Performance of RevPRAG on noisy datasets.
+Dataset Metrics GPT2-XL 1.5B Llama3-8B
+Noisy_NQ TPR 0.980 0.977
+FPR 0.034 0.012
+Noisy_HotpotQA TPR 0.969 0.974
+FPR 0.026 0.011
+Noisy_MS-MARCO TPR 0.989 0.977
+FPR 0.013 0.018
+5.6 Efficiency
+Table 8 compares the time overhead between LLM Factoscope He et al.
+(2024) and RevPRAG when the LLM in RAG is Llama3-8B, including the average training time per epoch and the average inference time per test sample.
+This experiment was conducted using 1,000 training samples and 500 test samples, with poisoned and clean examples each accounting for 50%.
+The results demonstrate that RevPRAG, with its task-specific architecture and carefully selected detection metrics, incurs significantly lower computational costs than LLM Factoscope, which integrates multiple sub-models for hallucination detection.
+Its efficient detection capability makes RevPRAG particularly well-suited for latency-sensitive RAG scenarios, underscoring its practical value.
+Table 8: Comparison of time overhead.
+Dataset Training Time per Epoch Inference Time per Sample
+LLM factoscope RevPRAG LLM factoscope RevPRAG
+NQ 91.61s 19.31s 0.0051s 0.0021s
+HotpotQA 101.25s 23.69s 0.0066s 0.0023s
+MS-MARCO 94.47s 20.72s 0.0058s 0.0022s
 6 Conclusion
@@
 Overall, we believe our approach can further promote the secure application of RAG.
-Appendix A Additional Results
-Table 6: Generalization Performance of RevPRAG for RAG with four different LLMs. The training and test datasets vary across different rows. Abbreviations: Hot (HotpotQA), MS (MS-MARCO).
-Training Dataset Test Dataset Metrics LLMs of RAG
-GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
-NQ & Hot MS TPR 0.881 0.886 0.948 0.956
-FPR 0.134 0.149 0.076 0.066
-Hot & MS NQ TPR 0.980 0.983 0.988 0.980
-FPR 0.007 0.074 0.078 0.038
-NQ & MS Hot TPR 0.977 0.961 0.942 0.978
-FPR 0.025 0.089 0.055 0.049
-NQ & Hot & MS NQ & Hot & MS TPR 0.986 0.994 0.985 0.987
-FPR 0.032 0.007 0.009 0.035
-A.1 Generalization
-Given the wide range of RAG application scenarios and the diverse user requirements it faces, it is impractical to ensure that our detection model has been trained on all possible scenarios and queries in real-world applications.
-However, the performance of neural network models largely depends on the similarity between the distributions of the training data and the test data Yang et al.
+Appendix A Training Details
+A.1 Dataset
+As shown in Table 9, we present the average response lengths for both poisoned and correct answers generated by GPT2-XL across three datasets (NQ, HotpotQA, and MS-MARCO), along with examples illustrating each answer format for a specific question.
+To evaluate the detection of poisoning attacks on the knowledge base of RAG, we selected 3,000 instances of triples $(q,t,a)$ from each of the three evaluation datasets mentioned above.
+In each triple, $q$ denotes a question, $t$ represents the supporting text collected from Wikipedia or web documents corresponding to $q$, and $a$ is the correct answer to $q$, generated using the state-of-the-art GPT-4 model.
+Among these 3,000 triplets, 1,500 are randomly selected as benign instances, while the remaining 1,500 are designated as poisoned instances.
+For each poisoned instance, the poisoned answer $a_{p}$ is generated by GPT-4 for the given question $q$, and the poisoned text $t_{p}$ is crafted using existing poisoning strategies, including PoisonedRAG Zou et al.
+(2024), GARAG Cho et al.
+(2024), and PRCAP Zhong et al.
 (2023).
-Consequently, our model’s performance may degrade when faced with training and test data that stem from differing distributions—a challenge frequently observed in real-world scenarios.
-To address this issue, we conduct a series of generalization experiments.
-Specifically, we train the detection model using any two datasets and test it on a third dataset that was not used during training.
-For example, we use NQ and HotpotQA as training datasets and MS-MARCO as the testing dataset.
-Although these three datasets are all QA datasets, they exhibit significant differences.
-For example, NQ focuses on extracting answers to factual questions from a single long document, HotpotQA involves multi-document reasoning to derive answers, and MS-MARCO retrieves and ranks relevant answers from a large-scale collection of documents.
-Therefore, conducting generalization experiments based on these three datasets is reasonable.
-Table 6 illustrates the TPRs and FPRs of RevPRAG for RAG with four different LLMs.
-Overall, the experimental results demonstrate that our detection model exhibits strong generalization performance across RAG with different LLMs and various datasets.
-For example, when using HotpotQA and MS-MARCO as training data, the detection model achieves TPRs of 98% (with GPT2-XL 1.5B), 98.3% (with Llama2-7B), 98.8% (with Mistral-7B), and 98% (with Llama3-8B) on the NQ dataset.
-Meanwhile, all FPRs remain below 8%.
-Furthermore, we observe that the generalization performance is best when NQ is used as the test data (for instance, 98.3% with Llama2-7B), while MS-MARCO shows the poorest performance (for instance, 88.6% with Llama2-7B).
-We attribute this to the fact that the questions and tasks in HotpotQA and MS-MARCO are more complex compared to those in NQ.
-Therefore, detection models trained on more complex tasks generalize well to simpler tasks, whereas the reverse is more challenging.
-In conclusion, these experimental results highlight that RevPRAG exhibits strong generalization and robust detection performance, even in the presence of significant discrepancies between the training and test datasets.
-(a) TPRs of RevPRAG.
-(b) FPRs of RevPRAG.
-Figure 7: RevPRAG trained on the activations from specific layers.
-A.2 RevPRAG’s Performance on Complex Open-Ended Questions
-In this section, we conducted a series of experiments to evaluate the performance of RevPRAG on complex, open-ended questions (e.g., “how to make relationship last?”).
-These questions present unique challenges due to their diverse and unstructured nature, in contrast to straightforward, closed-ended questions (e.g., “What is the name of the highest mountain?”).
-In our experiments, the NQ, HotpotQA, and MS-MARCO datasets primarily consist of close-ended questions.
-As a result, the majority of our previous experiments focused on close-ended problems, which was our default experimental setting.
-In this study, we utilized the advanced GPT-4o to filter and extract 3,000 open-ended questions from the HotpotQA and MS-MARCO datasets for training and testing the model.
-For open-ended questions, cosine similarity is employed to evaluate whether the LLM’s response aligns with the attacker’s target response.
-If the similarity surpasses a predefined threshold, it is considered indicative of a successful poisoning attack.
-The experimental results are shown in Table 7.
-We can observe that RevPRAG demonstrates excellent detection performance even on complex open-ended questions.
-For example, RevPRAG achieved TPRs of 99.1% on HotpotQA and 99.0% on MS-MARCO, alongside FPRs of 0.8% on HotpotQA and 0.1% on MS-MARCO for RAG utilizing the Mistral-7B model.
-Table 7: RevPRAG achieved high TPRs and low FPRs on the open-ended questions from HotpotQA and MS-MARCO datasets.
-Dataset Metrics LLMs of RAG
-GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
-HotpotQA TPR 0.982 0.995 0.991 0.982
-FPR 0.033 0.029 0.008 0.007
-MS-MARCO TPR 0.988 0.989 0.990 0.983
-FPR 0.009 0.009 0.001 0.017
-Table 8: RevPRAG trained on the activations from specific individual layers of GPT2-XL 1.5B.
-Dataset Metrics Different layers
-layer 0 layer 8 layer 15 layer 22 layer 29 layer 36 layer 41 layer 47
-NQ TPR 0.996 0.988 0.996 0.984 0.996 0.988 0.992 0.996
-FPR 0.027 0.007 0.017 0.003 0.007 0.003 0.017 0.003
-HotpotQA TPR 0.713 0.984 0.994 0.989 0.520 0.619 0.931 0.992
-FPR 0.409 0.023 0.012 0.006 0.445 0.409 0.023 0.019
-MS-MARCO TPR 0.967 0.998 0.988 0.986 0.988 0.963 0.955 0.992
-FPR 0.023 0.004 0.002 0.019 0.030 0.026 0.037 0.017
-A.3 Activations from Specified Layers.
-Fig.
-7 illustrates the detection performance of RevPRAG using activations from different layers of various LLMs.
-In previously presented experiments, we utilize activations from all layers as both training and testing data, yielding excellent results.
-Additionally, we also test using different layers.
-The experimental results in Fig.
-7 demonstrate that utilizing activations from only the first few layers can still achieve satisfactory detection performance, providing valuable insights for future research.
-For example, when using activations from layers 0 to 5, RevPRAG achieved TPRs exceeding 97% while maintaining FPRs below 7% for RAG with all LLMs on HotpotQA.
-However, the experimental results also suggest that using activations from intermediate or deeper layers can lead to performance fluctuations, including signs of degradation or slower convergence.
-For instance, when using activations from layers 16 to 24 with Llama3-8B as the LLM in RAG, RevPRAG achieves a TPR of 78.8% on NQ dataset and 86% on MS-MARCO dataset.
-We further explored the use of activations from a specific individual layer of the LLMs to train and test RevPRAG.
-We chose 8 layers with roughly even spacing for testing.
-As shown in Table 8, when using activations from only a specific layer of the GPT2-XL model, RevPRAG demonstrates excellent performance in general.
-For instance, when the model is trained using activations from layer 0 on the NQ dataset, the TPR can reach as high as 99.6%.
-However, we also observed that activations from certain layers do not yield satisfactory performance.
-For example, when the model is trained using activations from layer 29 on the HotpotQA dataset, the TPR is only 52%, while the FPR reaches 44.5%.
-It is precisely due to the existence of these suboptimal layers that models trained with multi-layer activations may not always outperform those using single-layer activations (such as layer 0 with NQ dataset).
-However, incorporating multi-layer activations can enhance the model’s stability, mitigating the detrimental effects of these suboptimal layers.
-A.4 Impact of Similarity Metric.
-Table 9 shows the results on the HotpotQA dataset, indicating that the choice of similarity calculation method has minimal impact on RevPRAG’s performance, which consistently achieves high TPR and low FPR.
-This suggests the robustness of our approach, as it reliably identifies poisoned texts even when LLM activations vary slightly under similar conditions.
-Table 9: Impact of similarity metric.
-Similarity Metric Metrics LLMs of RAG
-GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
-Dot Product TPR 0.972 0.985 0.977 0.973
-FPR 0.016 0.061 0.022 0.017
-Cosine TPR 0.978 0.990 0.979 0.981
-FPR 0.037 0.011 0.023 0.043
-A.5 Isolating Poisoned Responses and Hallucinations
-It is well-known that hallucinations are an inevitable phenomenon in LLMs.
-Even with the introduction of a knowledge database in RAG, LLMs may still generate non-factual responses due to hallucinations.
-Therefore, the incorrect responses generated by RAG may also stem from hallucinations, rather than being solely caused by RAG poisoning.
-We conducted experiments to test if our approach can distinguish hallucinations and RAG poisoning.
-Fig.
-8 shows the t-SNE representation of mean activations for poisoned response and hallucinations across all layers for Mistral-7B and Llama2-7B on the NQ dataset.
-We observe that activations across all layers clearly distinguish between hallucinations and poisoned responses.
-Figure 8: t-SNE visualizations of activations for poisoned responses and hallucinations.
-This key finding has led us to extend our approach to differentiate between poisoned responses and hallucinations.
-We thus continue to collect data and train the model using the process outlined in Fig.
-3, with the only difference being that we now collect hallucination data.
-We also conduct extensive experiments on RAG with different LLMs and datasets.
-From the experimental results in Table 10, we can see that our method achieves a high TPR across all LLMs and datasets.
-For instance, RevPRAG achieved 98.7% (on NQ), 97.5% (on HotpotQA), and 97.3% (on MS-MARCO) TPRs for RAG with GPT2-XL 1.5B.
-Furthermore, we observe that the FPR remains low across all evaluation settings.
-As shown in the table, RevPRAG could achieve 0.8% (on NQ), 0.8% (on HotpotQA) and 0.6% (on MS-MARCO) FPRs for RAG with Llama3-8B.
-This further supports our previous observation that there is a clear distinction between poisoned responses and hallucinations.
-Table 10: RevPRAG could achieve high TPRs and low FPRs to distinguish poisoned responses and hallucinations.
-Dataset Metrics LLMs of RAG
-GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
-NQ TPR 0.987 0.983 0.993 0.995
-FPR 0.046 0.017 0.069 0.008
-HotpotQA TPR 0.975 0.978 0.991 0.995
-FPR 0.004 0.058 0.004 0.008
-MS-MARCO TPR 0.973 0.984 0.999 0.989
-FPR 0.009 0.023 0.001 0.006
+The dataset is split into 70% for training, 20% for testing, and 10% as a support set.
+Within the training set, samples are randomly grouped into triplets (anchor, positive, negative), where the anchor and positive belong to the same class, and the negative belongs to a different class.
+Table 9: Statistical data and format of the responses.
+Dataset Average Word Count of Response An Example of Response
+NQ Poisoned Response: 7 Correct Response: 12 Question: where is the food stored in a yam plant? Poisoned Response: In the leaves. Correct Response: In the tuber.
+HotpotQA Poisoned Response: 8 Correct Response: 11 Question: Which actor starred in Assignment to Kill and passed away in 2000? Poisoned Response: Patrick O’Neal. Correct Response: John Gielgud.
+MS-MARCO Poisoned Response: 16 Correct Response: 24 Question: what is hardie plank?
+Poisoned Response: Hardie plank is a wood flooring option that is used for a variety of home styles.
+Correct Response: Hardie Plank is a brand of fiber cement siding.
+A.2 Poisoned Texts Generation
+To ensure that the retrieved poisoned texts successfully achieve the poisoning effect, we employ three existing methods PoisonedRAG Zou et al.
+(2024), GARAG Cho et al.
+(2024), and PRCAP Zhong et al.
+(2023) to generate the poisoned texts.
+In the PoisonedRAG Zou et al.
+(2024) method, the attacker first selects a target question along with its corresponding incorrect answer.
+The attacker then optimizes the design of the poisoned text to ensure that it meets two key criteria: (1) retrievability by the retriever and (2) effectiveness in misleading the language model to generate the incorrect answer.
+GARAG Cho et al.
+(2024) is a novel adversarial attack algorithm that generates adversarial documents by subtly perturbing clean ones while preserving answer tokens.
+Through iterative crossover, mutation, and selection, it optimizes the documents to maximize adversarial effectiveness within the defined search space.
+PRCAP Zhong et al.
+(2023) is a gradient-based method, which starts from a natural-language passage and iteratively perturbs it in the discrete token space to maximize its similarity to a set of training queries.
+It’s worth noting that the generation method for poisoned texts can be any approach that successfully achieves the poisoning effect.
+Once the activations of both correct and poisoned responses are obtained, we preprocess and use them for training and testing the RevPRAG model.
+A.3 Prompt
+The following is the system prompt for RAG, instructing an LLM to produce a response based on the provided context:
+Figure 6: The prompt used in RAG to make an LLM generate an answer based on the retrieved texts.
+A.4 Environment
+We conduct experiments on a server with 64 AMD EPYC 9654 CPUs (64 logical cores enabled) at 2.40–3.70 GHz, 512 GB of DDR5 RAM (assumed based on high-core-count server standards), and four NVIDIA RTX A6000 GPUs, each with 48 GB GDDR6 memory.